# Fixed-Point RNNs: Interpolating from Diagonal to Dense

### v2

@@ 4.2 FP-Mamba Iteration
Let us apply the the fixed-point RNN framework to the Mamba parametrization. We represent the hidden state as $\mathbf{h}_{t}^{\ell}$, where $t$ is the token index (i.e., indexing over the sequence dimension), and $\ell$ is the fixed-point iteration index (i.e., indexing over the depth dimension). The same notation is used for other variables to emphasize when they depend on the input and hidden state of the current iteration. We propose the following iteration to adapt Mamba to the fixed-point mechanism for matrix state RNNs in Eq. 11:
$\mathbf{h}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{h}_{t-1}^{\ell}+\mathbf{\bar{B}}_{t}^{\ell}\left(\Delta_{t}\mathbf{Q}_{t}^{\ell}x_{t}\right)^{\top}+\mathbf{\bar{B}}_{t}^{\ell}\left(\Delta_{t}\left(\mathbf{I}-\mathbf{Q}_{t}^{\ell}\right)y_{t}^{\ell-1}\right)^{\top},$ ${y_{t}^{\ell}}^{\top}=({\mathbf{\bar{C}}_{t}^{\ell}})^{\top}\mathbf{h}_{t}^{\ell}.$ (12)
L2-normalizing $\mathbf{\bar{B}}_{t}^{\ell}$ and $\mathbf{\bar{C}}_{t}^{\ell}$ allows to limit the Lipschitz constant according to Theorem 3.1. Furthermore, we replace the normalization term $(\mathbf{1}-\boldsymbol{\lambda}_{t})$ with Mamba’s normalization term $\Delta_{t}$. Expanding $y_{t}^{\ell-1}$ yields the recurrence on the matrix state
$\mathbf{h}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{h}_{t-1}^{\ell}+\mathbf{\bar{B}}_{t}^{\ell}(\Delta_{t}\mathbf{Q}_{t}^{\ell}x_{t})^{\top}+\mathbf{\bar{B}}_{t}^{\ell}(\mathbf{\bar{C}}_{t}^{\ell-1})^{\top}\mathbf{h}_{t}^{\ell-1}(\mathbf{I}-\mathbf{Q}_{t}^{\ell})^{\top}\Delta_{t},$
where the last term nicely illustrates the two components which mix the channels of the hidden states: the low-rank matrix $\mathbf{\bar{B}}_{t}^{\ell}(\mathbf{\bar{C}}_{t}^{\ell-1})^{\top}$ mixes over the dimension $d_{\text{state}}$, while $(\mathbf{I}-\mathbf{Q}_{t}^{\ell})^{\top}$ mixes over the dimension $d_{\text{inner}}$. This factorization significantly simplifies the fourth-order tensor mixer formulation introduced in Eq. 10 and performs well in practice.
Finally, Eq. 12 can be computed as Mamba with an adjusted input $\tilde{x}_{t}^{\ell}=\mathbf{Q}_{t}^{\ell}\left(x_{t}-y_{t}^{\ell-1}\right)+y_{t}^{\ell-1}$,
$\mathbf{h}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{h}_{t-1}^{\ell}+\mathbf{\bar{B}}_{t}^{\ell}\left(\Delta_{t}\tilde{x}_{t}^{\ell}\right)^{\top}.$ (13)
Dependence on $y_{t-1}^{\ell-1}$ Test Accuracy
@@
$0.11\quad\pm 0.00$
$0.53\quad\pm 0.02$
$0.45\quad\pm 0.05$
$0.55\quad\pm 0.05$
$0.81\quad\pm 0.01$
$0.88\quad\pm 0.01$
$0.86\quad\pm 0.02$
$0.94\quad\pm 0.03$
@@
Table 1: Effect of shifted hidden state dependence $y_{t-1}^{\ell-1}$ on copying at ×$\times 2$ length generalization. Each column determines which input-dependent component of the recurrence in Eq. 12 also depends on $y_{t-1}^{\ell-1}$. Performance is unlocked by including a hidden dependence for $\mathbf{B}_{t}$ and $\mathbf{C}_{t}$.
@@
Figure 5: Sequence length generalization on the copy task. A 1-layer FP-Mamba with Householder $\mathbf{Q}_{t}$ matches a 2-layer GatedDeltaNet baseline. Note that the median number of fixed-point iterations at test time (gray vertical line) is well below the longest train sequence length (pink line).
@@
In other words, one fixed-point step consists of a channel mixing using $\mathbf{Q}_{t}$, followed by a sequence mixing using Mamba. This separation of concerns allows to speed up the parallel recurrence in time using the Mamba implementation. To find a fixed-point, the two phases are repeated until convergence, i.e. $\frac{\|y^{\ell}-y^{\ell-1}\|_{\infty}}{\|y^{\ell}\|_{\infty}}<0.1$. After convergence to a fixed-point, $\mathbf{h}_{t}^{*}$ and $y_{t}^{*}$ present the hidden state and output of the dense matrix-valued RNN. Similar to Mamba, we apply a gated linear unit $\mathbf{G}_{t}\in\mathbb{R}^{d_{\text{inner}}}$ to the output, which we observe to provide a slight improvement in performance when present in the fixed-point loop: $\tilde{y}_{t}^{\ell}=\mathbf{G}_{t}\odot y_{t}^{\ell-1}$.
4.3 Shifted Hidden State Dependence $y_{t-1}^{\ell-1}$
@@
In preliminary experiments, we observe that even the Fixed-Point RNN with input-dependent parameters and matrix state akin to Mamba-1 is outperformed by Mamba-2 or DeltaNet (Dao & Gu, 2024; Yang et al., 2024b) on a copy task. Inspired by the short convolution in Mamba, we investigate the effect of augmenting the input-dependence of parameters $\boldsymbol{\lambda}_{t}^{\ell}$, $\mathbf{B}_{t}^{\ell}$, $\mathbf{C}_{t}^{\ell}$, and $\mathbf{Q}_{t}^{\ell}$ at iteration $\ell$ with a shifted hidden state dependence. In practice, this means that these are linear functions of $x_{t}$ as well as the shifted previous iterate in depth $y_{t-1}^{\ell-1}$. We refer the reader to App. C.2 for the exact formulation of the dependency.
@@
In Tab. 1, we ablate the hidden state dependence for various combinations of $\boldsymbol{\lambda}_{t}$, $\mathbf{B}_{t}$, $\mathbf{C}_{t}$, and a Householder $\mathbf{Q}_{t}$. Observe that the dependence of $\mathbf{B}_{t}$ and $\mathbf{C}_{t}$ is crucial to enable the model to copy. In App. C.4, we discuss why this dependence of $\mathbf{B}_{t}$ and $\mathbf{C}_{t}$ could be important for copying. If additionally $\boldsymbol{\lambda}_{t}$ and $\mathbf{Q}_{t}$ depend on $y_{t-1}^{\ell-1}$, the copy task is essentially solvable at ×$\times 2$ length generalization. We therefore adopt the hidden state dependence for all components in FP-Mamba.
@@
In Fig. 5, we evaluate length generalization on the copying task. While the best-performing baseline Gated DeltaNet is specifically designed for associative recall tasks (Yang et al., 2025), both Mamba 1 and 2 struggle with ×$\times 2$ generalization. FP-Mamba closes this gap and proves the effectiveness of our proposed modifications for better memory. We would like to highlight that the number of fixed-point iterations $\ell^{*}$ (gray vertical line) in FP-Mamba is well below the maximum sequence length.
@@
Figure 6: Length generalization as a function of training time on $A_{5}$. Wall clock time is plotted against the longest test sequence length with $>90\%$ accuracy for every model. While baselines of increasing depth cannot generalize beyond the training sequence length $16$ (horizontal pink line), our proposed framework allows to achieve much higher generalization by scaling training time through the number of fixed-point iterations $\ell_{\text{max}}$.
@@ 4.4 State-Tracking
Figure 7: The effective number of fixed-point iterations for each layer of a FP-Mamba throughout language pretraining on FineWeb (Penedo et al., 2024). Note that the number of fixed-point iterations $\ell^{*}$ stays well below the training context length of 2048. The corresponding validation perplexities are available in App. E.1.
4.5 Required Number of Iterations
A fixed-point iteration in the forward pass inevitably introduces sequential overhead to the computation of a model. While this might be acceptable for sequential generation at test time, reduced parallelism can be inhibiting at training time. In Fig. 1, we therefore evaluate FP-Mamba on $A_{5}$ with limited number of fixed-point iterations at training time $\ell_{\text{max}}\in\{2,4,8,16\}$. We observe that the performance decreases once $\ell_{\text{max}}$ is lower than the training sequence length of $16$. In Fig. 6, we confirm that the resulting longer training times are indeed required for good length generalization. However, as opposed to baselines of increasing depth $\in\{1,2,4,6,8\}$, fixed-point iterations are able to make use of the additional training time. Furthermore, there is room to improve efficiency, as suggested by a simple randomization scheme (gray stars) where $\ell_{\text{max}}\sim\Gamma(4,1)$ is sampled from a Gamma distribution with mean $4$ for every batch. But most importantly, the effective number of fixed-point iterations depends on the difficulty of the task. For example, in Fig. 7, we observe on language pretraining at context length 2048 that the model automatically adapts to using less fixed-point iterations presumably because the full state-tracking expressivity is not required.
@@ 5 Discussion
A fixed-point mechanism, such as the one introduced in this paper, endows a parallelizable, diagonal linear RNN with the ability to dynamically increase the sequential computation per token and describe a dense linear RNN in the limit. Our results show that such a paradigm can enable both strong state-tracking and memory capabilities with a constant number of parameters in a combined sequence and channel mixing layer (Fig. 2). In fact, the fixed-point iteration gradually transforms a diagonal (i.e., channel-wise) RNN into a dense (i.e., channel-mixing) RNN, thereby allowing to trade parallel computation for expressivity (Fig. 1) without incurring additional cost during backpropagation.
For Fixed-Point RNNs to become competitive in practice, it is paramount to further understand the trade-offs between parallel and sequential computation. In the worst case, if the sequential overhead is linear in the sequence length $T$, the model essentially behaves like a traditional, non-linear RNN but with quadratic runtime $O(T^{2})$. This, however, is not necessarily a disadvantage if the model is capable of adapting its sequential steps to the difficulty of the task, with negligible cost for less demanding tasks. In this paper, we focus on introducing and parameterizing the framework for Fixed-Point RNNs. Therefore, we leave a thorough analysis and improvement of fixed-point convergence speeds beyond our preliminary results (Fig. 13(c) 57) to future work.
Fixed-Point RNNs present an interesting opportunity to be fused into a single GPU kernel with reduced memory I/O. This is an inherent advantage from performing repeated computations on the same operands. Several open problems need to be solved to achieve that: (1) different implementations should converge to the same fixed-point, wether they compute the recurrence sequentially, in parallel, or in chunks, (2) the input-dependence and memory footprint of the fixed-point iteration should be reduced to satisfy current hardware limitations, and (3) alternative structures for sequence or channel mixers $\mathbf{Q}_{t}$ could reduce the computational burden. Future progress on these problems could enable significant speed-ups in practical implementations of Fixed-Point RNNs.
@@ Conclusion
We would like to thank Riccardo Gerazzi and Julien Siems for the helpful discussions and comments. Antonio Orvieto, Felix Sarnthein and Sajad Movahedi acknowledge the financial support of the Hector Foundation. Felix Sarnthein would also like to acknowledge the financial support from the Max Planck ETH Center for Learning Systems (CLS).
@@
The quadratic runtime complexity of Transformers motivated research on the linearization of its attention mechanism (Wang et al., 2020; Chen et al., 2021; Choromanski et al., 2020) – a technique that inevitably brings the sequence mixing mechanism closer to RNN-like processing (Katharopoulos et al., 2020; Schlag et al., 2021). Recently, improvements on the long-range-arena benchmark (Tay et al., 2020) with state-space models (Gu et al., 2022; Smith et al., 2023) sparked a renewed interest in recurrent models (Gu & Dao, 2023; Sun et al., 2023; De et al., 2024; Qin et al., 2024; Peng et al., 2024; Yang et al., 2024a). New efficient token mixing strategies such as Mamba (Gu & Dao, 2023) showcase impressive results in language modeling (Waleffe et al., 2024) while offering linear runtime complexity. These models are fundamentally diagonal linear RNNs, which enables parallel algorithms such as parallel scans (Martin & Cundy, 2018) and fast linear attention based implementations (Yang et al., 2024b; Dao & Gu, 2024).
@@
Several works propose toy tasks to identify specific shortcomings of modern architectures. Specifically, Beck et al. (2024) use the Chomsky hierarchy to organize formal language tasks, of which a modular arithmetic task remains unsolved. With similar motivations, Merrill & Sabharwal (2023a) introduce a set of word-problems for assessing state-tracking capabilities, among which the $A_{5}$ and $S_{5}$ tasks remain unsolved by Transformers and SSMs. Motivated by Transformers outperforming RNNs in memory capabilities, Jelassi et al. (2024) introduce a copying task as a fundamental benchmark for memory. We focus on these tasks to evaluate our Fixed-Point RNN framework.
@@
B.1 Proof for Theorem 3.1 (Lipschitz constant)
@@
$f_{\theta}(x,h)_{t}=\sum_{\tau=0}^{t}\mathbf{\Lambda}^{t-\tau}\left(\mathbf{I}-\mathbf{\Lambda}\right)\left(\mathbf{Q}\mathbf{B}_{\tau}x_{\tau}+\left(\mathbf{I}-\mathbf{Q}\right)h_{\tau}\right).$
@@
$\left\|f_{\theta}(x,h)_{t}-f_{\theta}(x,h^{\prime})_{t}\right\|_{2}<\left\|h-h^{\prime}\right\|_{2},$
where $h$ and $h^{\prime}$ are two arbitrary hidden states. From the unrolled form, this is equivalent to
$\left\|\sum_{\tau=0}^{t}\mathbf{\Lambda}^{t-\tau}\left(\mathbf{I}-\mathbf{\Lambda}\right)\left(\mathbf{I}-\mathbf{Q}\right)(h_{\tau}-h_{\tau}^{\prime})\right\|_{2}<\left\|h-h^{\prime}\right\|_{2}.$ (14)
@@
$\left\|\sum_{\tau=0}^{t}\mathbf{\Lambda}^{t-\tau}\left(\mathbf{I}-\mathbf{\Lambda}\right)\left(\mathbf{I}-\mathbf{Q}\right)(h_{\tau}-h_{\tau}^{\prime})\right\|_{2}\leq\left\|\sum_{\tau=0}^{t}\mathbf{\Lambda}^{t-\tau}\right\|_{2}\cdot\left\|\mathbf{I}-\mathbf{\Lambda}\right\|_{2}\cdot\left\|\mathbf{I}-\mathbf{Q}\right\|_{2}\cdot\left\|h_{\leq t}-h_{\leq t}^{\prime}\right\|_{2},$
where $h_{\leq t}$ corresponds to the concatenation of the hidden states $h_{\tau}$ for $\tau\leq t$. Now to prove this product is $<\left\|h-h^{\prime}\right\|_{2}$, consider the terms individually. Since $\left\|h_{\leq t}-h_{\leq t}^{\prime}\right\|_{2}\leq\left\|h-h^{\prime}\right\|_{2}$, the remaining terms need to be $<1$. Assuming $\mathbf{\Lambda}$ is contractive, we use the Neumann series $\sum_{\tau=0}^{t}\mathbf{\Lambda}^{t-\tau}\leq(\mathbf{I}-\mathbf{\Lambda})^{-1}$ and get

### v3

@@ 4.2 FP-Mamba Iteration
Let us apply the the fixed-point RNN framework to the Mamba parametrization. We represent the hidden state as $\mathbf{H}_{t}^{\ell}$, where $t$ is the token index (i.e., indexing over the sequence dimension), and $\ell$ is the fixed-point iteration index (i.e., indexing over the depth dimension). The same notation is used for other variables to emphasize when they depend on the input and hidden state of the current iteration. We propose the following iteration to adapt Mamba with notation from App. C.1 to the fixed-point mechanism for matrix state RNNs in Eq. 11:
$\mathbf{H}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{H}_{t-1}^{\ell}+\mathbf{\bar{b}}_{t}^{\ell}\left(\Delta_{t}\mathbf{Q}_{t}^{\ell}\mathbf{x}_{t}\right)^{\top}+\mathbf{\bar{b}}_{t}^{\ell}\left(\Delta_{t}\left(\mathbf{I}-\mathbf{Q}_{t}^{\ell}\right)\mathbf{y}_{t}^{\ell-1}\right)^{\top},$ ${\mathbf{y}_{t}^{\ell}}^{\top}=({\mathbf{\bar{c}}_{t}^{\ell}})^{\top}\mathbf{H}_{t}^{\ell}.$ (12)
L2-normalizing $\mathbf{\bar{b}}_{t}^{\ell}$ and $\mathbf{\bar{c}}_{t}^{\ell}$ allows to limit the Lipschitz constant according to Theorem 3.1. Furthermore, we replace the normalization term $(\mathbf{1}-\boldsymbol{\lambda}_{t})$ with Mamba’s normalization term $\Delta_{t}$. Expanding $\mathbf{y}_{t}^{\ell-1}$ yields the recurrence on the matrix state
$\mathbf{H}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{H}_{t-1}^{\ell}+\mathbf{\bar{b}}_{t}^{\ell}(\Delta_{t}\mathbf{Q}_{t}^{\ell}\mathbf{x}_{t})^{\top}+\mathbf{\bar{b}}_{t}^{\ell}(\mathbf{\bar{c}}_{t}^{\ell-1})^{\top}\mathbf{H}_{t}^{\ell-1}(\mathbf{I}-\mathbf{Q}_{t}^{\ell})^{\top}\Delta_{t},$ (13)
where the last term nicely illustrates the two components which mix the channels of the hidden states: the low-rank matrix $\mathbf{\bar{b}}_{t}^{\ell}(\mathbf{\bar{c}}_{t}^{\ell-1})^{\top}$ mixes over the dimension $d_{\text{state}}$, while $(\mathbf{I}-\mathbf{Q}_{t}^{\ell})^{\top}$ mixes over the dimension $d_{\text{inner}}$. This factorization significantly simplifies the fourth-order tensor mixer formulation introduced in Eq. 10, remains expressive as discussed in App. F, and performs well in practice.
Finally, Eq. 12 can be computed as Mamba with an adjusted input $\mathbf{\tilde{x}}_{t}^{\ell}=\mathbf{Q}_{t}^{\ell}\left(\mathbf{x}_{t}-\mathbf{y}_{t}^{\ell-1}\right)+\mathbf{y}_{t}^{\ell-1}$,
$\mathbf{H}_{t}^{\ell}=\boldsymbol{\lambda}_{t}\odot\mathbf{H}_{t-1}^{\ell}+\mathbf{\bar{b}}_{t}^{\ell}\left(\Delta_{t}\mathbf{\tilde{x}}_{t}^{\ell}\right)^{\top}.$ (14)
Dependence on $\mathbf{y}_{t-1}^{\ell-1}$ Test Accuracy
@@
$0.11\qquad\pm 0.00$
$0.53\qquad\pm 0.02$
$0.45\qquad\pm 0.05$
$0.55\qquad\pm 0.05$
$0.81\qquad\pm 0.01$
$0.88\qquad\pm 0.01$
$0.86\qquad\pm 0.02$
$0.94\qquad\pm 0.03$
@@
Table 1: Effect of shifted hidden state dependence $\mathbf{y}_{t-1}^{\ell-1}$ on copying at $\times 2$ length generalization. Each column determines which input-dependent component of the recurrence in Eq. 12 also depends on $\mathbf{y}_{t-1}^{\ell-1}$. Performance is unlocked by including a hidden dependence for $\mathbf{b}_{t}$ and $\mathbf{c}_{t}$.
@@
Figure 5: Sequence length generalization on the copy task. A 1-layer FP-Mamba-H matches a 2-layer GatedDeltaNet baseline. Note that the median number of fixed-point iterations at test time $\ell^{*}$ (gray vertical line) is well below the longest training sequence length (pink line).
@@
In other words, one fixed-point step consists of a channel mixing using $\mathbf{Q}_{t}$, followed by a sequence mixing using Mamba. This separation of concerns allows to speed up the parallel recurrence in time using the Mamba implementation. To find a fixed-point, the two phases are repeated until $\frac{\|\mathbf{y}^{\ell}-\mathbf{y}^{\ell-1}\|_{\infty}}{\|\mathbf{y}^{\ell}\|_{\infty}}<0.1$ is satisfied. After these $\ell^{*}$ iterations, required for the model to converge to a fixed-point, $\mathbf{H}_{t}^{*}$ and $\mathbf{y}_{t}^{*}$ present the hidden state and output of the dense matrix-valued RNN $F_{\boldsymbol{\theta}}$. Similar to Mamba, we apply a gated linear unit $\mathbf{g}_{t}\in\mathbb{R}^{d_{\text{inner}}}$ to the output, which we observe to provide a slight improvement in performance when present within the fixed-point loop: $\mathbf{\tilde{y}}_{t}^{\ell}=\mathbf{g}_{t}\odot\mathbf{y}_{t}^{\ell-1}$.
4.3 Shifted Hidden State Dependence $\mathbf{y}_{t-1}^{\ell-1}$
@@
In preliminary experiments, we observe that even the Fixed-Point RNN with input-dependent parameters and matrix state akin to Mamba-1 is outperformed by Mamba-2 or DeltaNet (Dao & Gu, 2024; Yang et al., 2024b) on a copy task. Inspired by the short convolution in Mamba, we investigate the effect of augmenting the input-dependence of parameters $\boldsymbol{\lambda}_{t}^{\ell}$, $\mathbf{b}_{t}^{\ell}$, $\mathbf{c}_{t}^{\ell}$, and $\mathbf{Q}_{t}^{\ell}$ at iteration $\ell$ with a shifted hidden state dependence. In practice, this means that these are linear functions of $\mathbf{x}_{t}$ as well as the shifted previous iterate in depth $\mathbf{y}_{t-1}^{\ell-1}$. We refer the reader to App. C.2 for the exact formulation of the dependency.
@@
In Tab. 1, we ablate the hidden state dependence for various combinations of $\boldsymbol{\lambda}_{t}$, $\mathbf{b}_{t}$, $\mathbf{c}_{t}$, and a Householder $\mathbf{Q}_{t}$. Observe that the dependence of $\mathbf{b}_{t}$ and $\mathbf{c}_{t}$ is crucial to enable the model to copy. In App. C.4, we discuss why this dependence of $\mathbf{b}_{t}$ and $\mathbf{c}_{t}$ could be important for copying. If additionally $\boldsymbol{\lambda}_{t}$ and $\mathbf{Q}_{t}$ depend on $\mathbf{y}_{t-1}^{\ell-1}$, the copy task is essentially solvable at $\times 2$ length generalization. We therefore adopt the hidden state dependence for all components in FP-Mamba.
@@
In Fig. 5, we evaluate length generalization on the copying task. While the best-performing baseline Gated DeltaNet is specifically designed for associative recall tasks (Yang et al., 2025), both Mamba 1 and 2 struggle with $\times 2$ generalization. FP-Mamba closes this gap and proves the effectiveness of our proposed modifications for better memory. We would like to highlight that the number of fixed-point iterations $\ell^{*}$ (gray vertical line) in FP-Mamba is well below the maximum sequence length.
@@
Figure 6: Length generalization as a function of training time on $A_{5}$. Wall clock time is plotted against the longest test sequence length with $90\%$ accuracy for every model. While baselines of increasing depth cannot generalize beyond the training sequence length $16$ (horizontal pink line), our proposed framework allows to achieve much higher generalization by scaling training time through the number of fixed-point iterations $\ell$.
@@ 4.4 State-Tracking
Figure 7: The effective number of fixed-point iterations for each layer of a FP-Mamba-H throughout language pretraining on FineWeb (Penedo et al., 2024) at context length 2048. The corresponding validation perplexities are available in App. E.1.
4.5 Required Number of Iterations $\ell^{*}$
A fixed-point iteration in the forward pass inevitably introduces sequential overhead to the computation of a model. While this might be acceptable for sequential generation at test time, reduced parallelism can be inhibiting at training time. In Fig. 1, we therefore evaluate FP-Mamba-H on $A_{5}$ with limited number of fixed-point iterations at training time $\ell_{\text{max}}\in\{2,4,8,16\}$. We observe that the performance decreases once $\ell_{\text{max}}$ is lower than the training sequence length of $16$. In Fig. 6, we confirm that the resulting longer training times are indeed required for good length generalization. However, as opposed to baselines of increasing depth $\in\{1,2,4,6,8\}$, fixed-point iterations gain from the additional training time. Furthermore, there is room to improve efficiency, as suggested by a simple randomization scheme (gray stars) where $\ell_{\text{max}}\sim\Gamma(4,1)$ is sampled from a Gamma distribution with mean $4$ for every batch. But most importantly, the effective number of fixed-point iterations depends on the difficulty of the task. Indeed, Fig. 7 shows that the model automatically adapts to using less fixed-point iterations on language pretraining at context length 2048. Similarly, on copying (Fig. 5) and and modular arithmetic (Fig. 10), we observe that the required number of fixed-point iterations $\ell^{*}$ is well below the sequence length $T$. This suggests that the model adapts to $O(T)$ complexity on simpler tasks when the full state-tracking expressivity is not required.
Forward Backward
Mamba $O(T)$ $O(T)$
FP-Mamba $O\left((T+C_{\mathbf{Q}_{t}})\cdot\min(\ell^{*},\ell_{\text{max}})\right)$ $O(T+C_{\mathbf{Q}_{t}})$
Table 2: Complexity of FP-Mamba in comparison to Mamba. The cost of channel mixing with structure $\mathbf{Q}_{t}$ is denoted by $C_{\mathbf{Q}_{t}}$.
@@ 5 Discussion
A fixed-point mechanism, such as the one introduced in this paper, endows a parallelizable, diagonal linear RNN with the ability to dynamically increase the sequential computation and describe a dense linear RNN in the limit. Our results show that such a paradigm can enable both strong state-tracking and memory capabilities with a constant number of parameters in a combined sequence and channel mixing layer (Fig. 2). In fact, the fixed-point iteration gradually transforms a diagonal (i.e., channel-wise) RNN into a dense (i.e., channel-mixing) RNN, thereby allowing to trade parallel computation for expressivity (Fig. 1) without incurring additional cost during backpropagation (cf. Tab. 2).
For Fixed-Point RNNs to become competitive in practice, it is important to further understand the trade-offs between parallel and sequential computation. In the worst case, as shown in Tab. 2, FP-RNNs could behave like traditional, non-linear RNNs with quadratic runtime $O(T^{2})$ if the sequential overhead $\ell^{*}$ is linear in the sequence length $T$. This, however, is not necessarily a disadvantage since FP-RNNs adapt $\ell^{*}$ to the difficulty of the task. In this paper, we focus on introducing the framework for FP-RNNs and leave the improvement of fixed-point convergence rates to future work.
Fixed-Point RNNs present an interesting opportunity to be fused into a single GPU kernel with reduced memory I/O. This is an inherent advantage from performing repeated computation on the same operands. Several open problems need to be solved to achieve that: (1) different implementations such as sequential, parallel, or chunk-wise should converge to the same fixed-points, (2) the memory footprint of the fixed-point iteration should satisfy current hardware limitations, and (3) alternative sequence or channel mixer structures could unlock higher efficiency. Future progress on these problems could enable significant speed-ups in practical implementations of Fixed-Point RNNs.
@@ Conclusion
We would like to thank Riccardo Grazzi and Julien Siems for the helpful discussions and comments. Antonio Orvieto, Felix Sarnthein and Sajad Movahedi acknowledge the financial support of the Hector Foundation. Felix Sarnthein would also like to acknowledge the financial support from the Max Planck ETH Center for Learning Systems (CLS).
@@
The quadratic runtime complexity of Transformers motivated research on the linearization of its attention mechanism (Wang et al., 2020; Chen et al., 2021; Choromanski et al., 2020) – a technique that inevitably brings the sequence mixing mechanism closer to RNN-like processing (Katharopoulos et al., 2020; Schlag et al., 2021a). Recently, improvements on the long-range-arena benchmark (Tay et al., 2020) with state-space models (Gu et al., 2022; Smith et al., 2023) sparked a renewed interest in recurrent models (Gu & Dao, 2024; Sun et al., 2023; De et al., 2024; Qin et al., 2024; Peng et al., 2024; Yang et al., 2024a). New efficient token mixing strategies such as Mamba (Gu & Dao, 2024) showcase impressive results in language modeling (Waleffe et al., 2024) while offering linear runtime complexity. These models are fundamentally diagonal linear RNNs, which enables parallel algorithms such as parallel scans (Martin & Cundy, 2018) and fast linear attention based implementations (Yang et al., 2024b; Dao & Gu, 2024).
@@
Several works propose toy tasks to identify specific shortcomings of modern architectures. Specifically, Beck et al. (2024) use the Chomsky hierarchy to organize formal language tasks, of which a modular arithmetic task remains unsolved. With similar motivations, Merrill & Sabharwal (2023) introduce a set of word-problems for assessing state-tracking capabilities, among which the $A_{5}$ and $S_{5}$ tasks remain unsolved by Transformers and SSMs. Motivated by Transformers outperforming RNNs in memory capabilities, Jelassi et al. (2024) introduce a copying task as a fundamental benchmark for memory. We focus on these tasks to evaluate our Fixed-Point RNN framework.
@@
B.1 Proof for Theorem 3.1 (Lipschitz constant of $f_{\theta}(\mathbf{x},\mathbf{h})$ is <1<1)
See 3.1
@@
$f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})_{t}=\sum_{\tau=0}^{t}\mathbf{\Lambda}^{t-\tau}\left(\mathbf{I}-\mathbf{\Lambda}\right)\left(\mathbf{Q}\mathbf{B}_{\tau}\mathbf{x}_{\tau}+\left(\mathbf{I}-\mathbf{Q}\right)\mathbf{h}_{\tau}\right).$
@@
$\left\|f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})_{t}-f_{{\boldsymbol{\theta}}}(\mathbf{x},\mathbf{h}^{\prime})_{t}\right\|_{2}<\left\|\mathbf{h}-\mathbf{h}^{\prime}\right\|_{2},$
where $\mathbf{h}$ and $\mathbf{h}^{\prime}$ are two arbitrary hidden states. From the unrolled form, this is equivalent to
$\left\|\sum_{\tau=0}^{t}\mathbf{\Lambda}^{t-\tau}\left(\mathbf{I}-\mathbf{\Lambda}\right)\left(\mathbf{I}-\mathbf{Q}\right)(\mathbf{h}_{\tau}-\mathbf{h}_{\tau}^{\prime})\right\|_{2}<\left\|\mathbf{h}-\mathbf{h}^{\prime}\right\|_{2}.$ (15)
@@
$\left\|\sum_{\tau=0}^{t}\mathbf{\Lambda}^{t-\tau}\left(\mathbf{I}-\mathbf{\Lambda}\right)\left(\mathbf{I}-\mathbf{Q}\right)(\mathbf{h}_{\tau}-\mathbf{h}_{\tau}^{\prime})\right\|_{2}\leq\left\|\sum_{\tau=0}^{t}\mathbf{\Lambda}^{t-\tau}\right\|_{2}\cdot\left\|\mathbf{I}-\mathbf{\Lambda}\right\|_{2}\cdot\left\|\mathbf{I}-\mathbf{Q}\right\|_{2}\cdot\left\|\mathbf{h}_{\leq t}-\mathbf{h}_{\leq t}^{\prime}\right\|_{2},$
where $\mathbf{h}_{\leq t}$ corresponds to the concatenation of the hidden states $\mathbf{h}_{\tau}$ for $\tau\leq t$. Now to prove this product is $\left\|\mathbf{h}-\mathbf{h}^{\prime}\right\|_{2}$, consider the terms individually. Since $\left\|\mathbf{h}_{\leq t}-\mathbf{h}_{\leq t}^{\prime}\right\|_{2}\leq\left\|\mathbf{h}-\mathbf{h}^{\prime}\right\|_{2}$, the remaining terms need to be <1<1. Assuming $\mathbf{\Lambda}$ is contractive, we use the Neumann series $\sum_{\tau=0}^{t}\mathbf{\Lambda}^{t-\tau}\leq(\mathbf{I}-\mathbf{\Lambda})^{-1}$ and get